# Astute RAG: Overcoming Imperfect Retrieval Augmentation and Knowledge Conflicts for Large Language Models

### v1

@@ 6 Conclusion
Our paper investigates the impact of imperfect retrieval on the performance of RAG systems and identifies knowledge conflicts as a key challenge. To address this, we introduce Astute RAG, a novel approach that leverages the internal knowledge of LLMs and iteratively refines the generated responses by consolidating internal and external knowledge in a source way. Our empirical results demonstrate the effectiveness of Astute RAG in mitigating the negative effects of imperfect retrieval and improving the robustness of RAG systems, particularly in challenging scenarios with unreliable external sources.
Among the limitations, Astute RAG’s effectiveness hinges on the capabilities of advanced LLMs with strong instruction-following and reasoning abilities, hence potentially more limited applicability with less sophisticated LLMs. As an important future direction, extending the experimental setup to include longer outputs would be important, where the challenges of imperfect retrieval and knowledge conflicts may be even more pronounced. Furthermore, a comprehensive analysis of the impact of various context types (Balachandran et al., 2024) would enhance the understanding of the proposed method’s effectiveness. Future work can also extend our method beyond LLMs and RAG, such as addressing knowledge conflicts in multimodal settings (Zhu et al., 2024).
@@ Acknowledgement
[…]
@@ Appendix B Data Collection
Encompassing a diverse range of natural questions, our benchmark consists of realistic retrieval results with Google Search [^9] https://developers.google.com/custom-search/v1/overview as the retriever and the Web as the corpus. Notably, we do not select questions or annotate answers based on the retrieval results. This setting allows us to analyze the severity of imperfect retrieval in real-world RAG. It distinguishes our benchmark from previous ones that employ synthetic retrieval corruptions or that unintentionally reduce the frequency of imperfect retrieval with biased construction protocols (Chen et al., 2024a; Yang et al., 2024). Overall, our benchmark contains 1,042 short-form question-answer pairs, each paired with 10 retrieved passages.
@@
Question-answer pairs. We consider question-answer pairs from four datasets of different properties spanning across general questions, domain-specific questions, and long-tail questions. NQ (Kwiatkowski et al., 2019) and TriviaQA (Joshi et al., 2017) are two widely-studied question-answering (QA) datasets in general domains. BioASQ (Tsatsaronis et al., 2015) is from biomedical domain that has demonstrated significant benefits from RAG when general-purpose LLMs are considered. PopQA (Mallen et al., 2023) focuses on long-tail knowledge and has been shown to be challenging for even advanced LLMs to solve without external knowledge. All these datasets contain questions with short-form answers and most of them list all valid answer variants. This format can support automatic verification of answer appearance in retrieved passages and model responses, leading to more precise evaluations.
@@
Retrieval process. For each question in our benchmark, we query Google Search to retrieve the top 30 results and select the first 10 accessible websites. From each retrieved website, we extract the paragraph corresponding to the snippet provided in Google Search results as the retrieved passage. We do not consider enhancements to the retrieval side, such as query rewriting, as such enhancements are typically already incorporated into commercial information retrieval systems.

### v2

@@ 6 Conclusion
We investigate the impact of imperfect retrieval on the performance of RAG systems and identify knowledge conflicts as a key challenge. To address this, we introduce Astute RAG, a novel approach that leverages the internal knowledge of LLMs and iteratively refines the generated responses by consolidating internal and external knowledge in a source way. We demonstrate the effectiveness of Astute RAG in mitigating the negative effects of imperfect retrieval and improving the robustness of RAG, particularly in challenging scenarios with unreliable external sources.
@@ Acknowledgement
Limitations
Astute RAG’s effectiveness hinges on the capabilities of advanced LLMs with strong instruction-following and reasoning abilities, hence potentially more limited applicability with less sophisticated LLMs. As an important future direction, extending the experimental setup to include longer inputs would be important, where the challenges of imperfect retrieval and knowledge conflicts may be even more pronounced.
@@ Appendix A Prompt Template for Astute RAG
Prompt for Intermediate Step Evaluation **Task:** You are provided with the following: 1. A question. 2. The correct answer. 3. The input context. 4. The model’s response, which contains: - Consolidated context. - Confidence scores for candidate answers. Your task is to: - Evaluate the **quality of the consolidated context** in the model’s response and provide a label: ‘<<<consolidation>>> correct <<</consolidation>>>’ or ‘<<<consolidation>>> incorrect <<</consolidation>>>’. This evaluation is only about whether the consolidation is correct given the input context. - Evaluate the **accuracy of the confidence score** (whether it aligns with the confidence of the supporting context) and provide a label: ‘<<<confidence>>> correct <<</confidence>>>’ or ‘<<<confidence>>> incorrect <<</confidence>>>’. The evaluation is only based on the consolidated context. Note that correct consolidation and confidence do not necessarily indicate the correct answer. Question: {query} Correct Answer: {answer} Input Context: {input} Model Response: {response} Evaluation:
@@ Appendix B Data Collection
Encompassing a diverse range of natural questions, our benchmark consists of realistic retrieval results with Google Search [^8] https://developers.google.com/custom-search/v1/overview as the retriever and the Web as the corpus. Notably, we do not select questions or annotate answers based on the retrieval results. This setting allows us to analyze the severity of imperfect retrieval in real-world RAG. It distinguishes our benchmark from previous ones that employ synthetic retrieval corruptions or that unintentionally reduce the frequency of imperfect retrieval with biased construction protocols (Chen et al., 2024a; Yang et al., 2024). Overall, our benchmark contains 1,042 short-form question-answer pairs, each paired with 10 retrieved passages. When collecting the passages, we retrieve the top 30 results and select the first 10 accessible websites. From each retrieved website, we extract the paragraph corresponding to the snippet provided in the search results as the retrieved passage. Retrieved results might contain natural noise with irrelevant or misleading information. We do not consider enhancements to the retrieval side, such as query rewriting, as such enhancements are typically already incorporated into commercial information retrieval systems. All of these datasets are short-form QA. Following previous work (Xiang et al., 2024; Wei et al., 2024; Mallen et al., 2023), a model response is considered correct if it contains the ground-truth answer. To enhance evaluation reliability, we prompt LLMs to enclose the exact answer within special tokens, extracting them as the final responses.
Question-answer pairs.
@@
We consider question-answer pairs from four datasets of different properties spanning across general questions, domain-specific questions, and long-tail questions. NQ (Kwiatkowski et al., 2019) and TriviaQA (Joshi et al., 2017) are two widely-studied question-answering (QA) datasets in general domains. BioASQ (Tsatsaronis et al., 2015) is from biomedical domain that has demonstrated significant benefits from RAG when general-purpose LLMs are considered. PopQA (Mallen et al., 2023) focuses on long-tail knowledge and has been shown to be challenging for even advanced LLMs to solve without external knowledge. All these datasets contain questions with short-form answers and most of them list all valid answer variants. This format can support automatic verification of answer appearance in retrieved passages and model responses, leading to more precise evaluations.
Retrieval process.
@@
For each question in our benchmark, we query Google Search to retrieve the top 30 results and select the first 10 accessible websites. From each retrieved website, we extract the paragraph corresponding to the snippet provided in Google Search results as the retrieved passage. We do not consider enhancements to the retrieval side, such as query rewriting, as such enhancements are typically already incorporated into commercial information retrieval systems.
Appendix C Comparison with Answer Refinement
In Tab. 4, we further compare Astute RAG with Answer Refinement, where the LLM is prompted to refine its initial answer by reconsidering external context. Notably, this baseline performs nearly identically to Self-Route, which is expected, because both approaches rely on the model itself to determine whether to revise its initial answer based on external knowledge. The comparison underscores that simply enabling LLMs to self-correct does not yield significant improvements beyond existing routing strategies. In contrast, Astute RAG continues to outperform all baselines across datasets, reinforcing the benefit of explicitly structured consolidation mechanisms.
Model NQ TriviaQA BioASQ PopQA Overall
Self-Route 47.5 79.9 58.0 38.2 57.6
Answer Refinement 47.1 79.9 58.0 38.2 57.5
Astute RAG 50.2 81.6 58.0 40.5 59.2
Table 4: Comparison of Astute RAG with routing and refinement-based baselines on Gemini.
Appendix D Comparison with Context Filtering
To isolate the effect of identifying irrelevant information separate from consolidation, we conducted an ablation study introducing two Context Filtering baselines. The first baseline filters only the retrieved documents, while the second filters both retrieved and generated documents prior to answer generation, both without applying the consolidation step. As shown in Tab. 5, while context filtering improves performance over the basic RAG baseline, it falls short of the performance achieved by Astute RAG. This supports our hypothesis that consolidating diverse information, including consistent, conflicting, and relevant content, rather than merely filtering out irrelevant parts, is critical to the performance gains observed with Astute RAG.
Model NQ TriviaQA BioASQ PopQA Overall
RAG 42.7 76.0 55.2 33.7 53.7
Context Filtering (retrieved) 43.7 77.0 57.0 34.3 54.8
Context Filtering (generated and retrieved) 49.2 79.2 56.6 40.4 57.9
Astute RAG 50.2 81.6 58.0 40.5 59.2
Table 5: Comparison with context filtering baselines.
Appendix E Comparison with Context Compression
Context compression Wang et al. (2023b); Yoon et al. (2024) is also a related direction. We further conduct experiments comparing our method with CompAct Yoon et al. (2024). The results in Tab. 6 and Tab. 7 show that context compression is ineffective in handling the challenges of imperfect context and knowledge conflicts. Notably, it even performs worse than the No RAG and RAG baselines in this context. The primary reason for this underperformance lies in the limitations of context compression. It struggles to effectively identify incorrect information when there are conflicts in context and often removes or reduces the appearance of helpful information in the process. This reinforces the importance of our approach, which does not rely solely on compression but instead integrates both internal and external knowledge while handling conflicts in a more nuanced manner.
Method NQ TriviaQA BioASQ PopQA Overall
No RAG 47.1 82.0 50.4 29.8 54.5
RAG 44.4 76.7 58.0 36.0 55.5
CompAct 38.6 68.9 49.3 30.3 48.4
Astute RAG 52.2 84.1 60.1 44.4 61.7
Table 6: Comparison with context compression on Claude.
Method NQ TriviaQA BioASQ PopQA Overall
No RAG 44.8 80.2 45.8 25.3 51.3
RAG 42.7 76.0 55.2 33.7 53.7
CompAct 35.3 65.0 47.6 30.9 46.0
Astute RAG 50.2 81.6 58.0 40.5 59.2
Table 7: Comparison with context compression on Gemini.
Appendix F Influence of passage ordering.
We apply different ordering strategies (Alessio et al., 2024) on RAG and Astute RAG. As shown in Tab. 8, we find that the improvement with Astute RAG is significantly larger than the gap between different ordering strategies. Moreover, the consolidation process makes Astute RAG less sensitive to it.
Method Ordering Strategy NQ TriviaQA BioASQ PopQA Overall
RAG Random 43.39 76.33 56.99 34.83 54.61
Ascending 43.05 75.62 57.69 34.83 54.51
Descending 44.41 76.68 58.04 35.96 55.47
Ping-pong Descending Top-to-bottom 44.75 77.39 57.69 35.96 55.66
Ping-pong Descending Bottom-to-top 44.41 75.62 58.04 35.96 55.18
AstuteRAG Random 51.86 84.81 61.19 41.57 61.61
Ascending 51.86 85.51 59.79 42.13 61.52
Descending 52.20 84.10 60.14 44.38 61.71
Ping-pong Descending Top-to-bottom 52.20 84.45 59.09 43.82 61.42
Ping-pong Descending Bottom-to-top 51.19 85.16 61.54 43.82 62.00
Table 8: Performance by Ordering Strategies.